Can you put this equation in Standard Form? Y=(-8,0), M=4
Answer by Jolliano(16)

Using the formula for one point form.
(y - y1) = m(x - x1).
Substituting your values in the equation above.
(y - 0) = 4(x - [-8])
y = 4(x + 8)
y = 4x + 32